OpenStudy (anonymous):

(4x+1)(4x+1)..... multiply.... I tried several times and don't get it

OpenStudy (anonymous):

I followed the book's directions and all it did was confuse me

OpenStudy (anonymous):

you use the formula (a+b)(a+b) = (axa) +(axb)+(bxa)+(bxb) = a^2 + ab+ba +b^2 =a^2+b^2

OpenStudy (anonymous):

so for you question all you do is the same process (4x x 4x)+(4x x 1) +(1 x 4x) +(1 x 1)= 16x^2 +4x+4x+1 =16x^2 +8x +1

OpenStudy (anonymous):

multiply everything by everything :D its just distribution, i can draw you a nice picture if you want that makes it make sense

OpenStudy (anonymous):

so I write it out like this.... (4x4)+(4x1)+(1x4)+1x1) ???

OpenStudy (anonymous):

dont leave out your Xs...

OpenStudy (anonymous):

(4x x 4x)+(4x x 1) +(1 x 4x) +(1 x 1)= 16x^2 +4x+4x+1 =16x^2 +8x +1

OpenStudy (anonymous):

old school F.O.I.L. approach : ) F)irst O)uter I)nner L)ast it may help when you're first learning distributive (i.e. multiplying each figure in the first equation with each in the second and then combining like terms)

jhonyy9 (jhonyy9):

you dont know formula (a+b)^2 =a^2 +2ab +b^2 ? - so you need using this ...